# Sequential model confidence sets

### v1

@@ 5.2 Postprocessing of wind gust predictions
Currently, weather services use numerical weather prediction (NWP) models for predicting the future state of the atmosphere. These NWP models try to quantify the forecast uncertainty using ensemble predictions, where each ensemble member represents a different future scenario. However, the resulting ensemble forecasts fail to adequately quantify the forecast uncertainty, as they are subject to systematic biases and dispersion errors. Hence, the forecasts ought to be statistically postprocessed to generate accurate and reliable predictions (Gneiting and Raftery, 2005; Vannitsem et al., 2018).
Schulz and Lerch (2022) present a comprehensive comparison of multiple approaches for statistical postprocessing of ensemble forecasts for wind gusts. Here, we will build on their case study and compare the different forecasting models for probabilistic wind gust prediction on an extended time period by using SMCSs under the weak hypothesis.
@@ 5.2.1 Data
In this application, we use the data set of Primo et al. (2024), which is an extension of the original data in Schulz and Lerch (2022). The forecast data is based on an operational NWP model from the German weather service (Deutscher Wetterdienst, DWD) in the period from 08/12/2010 to 30/06/2023 with corresponding observations at 174 geographically diverse weather stations in Germany. The forecasts are initialized at 00 UTC and have a forecast horizon of up to 21 hours. Here, we consider only those forecasts with a lead time of 18h referring to 18 UTC and the $K=166$ stations with at least 2,600 observations.
Next to the ensemble forecasts (EPS), we include the eight postprocessing methods applied by Schulz and Lerch (2022) in our comparison. Statistical postprocessing methods are typically distributional regression models that use ensemble predictions from the NWP model as input data. The methods in our comparison can be divided into three groups: Basic techniques that only make use of the underlying wind gust ensemble predictions (EMOS, MBM, IDR), established machine learning methods for postprocessing that incorporate additional covariates (EMOS-GB, QRF), and neural network-based approaches (DRN, BQN, HEN). While the first two groups fit a separate model for each station, the neural networks estimate one locally adaptive model for all stations. Further, the models differ in the resulting forecast distributions, which range from parametric (EMOS, EMOS-GB, DRN) to semi- and non-parametric types (all others). The postprocessing models are trained on the period from 2010 to 2015 using the same configuration as in Schulz and Lerch (2022) and evaluated on the remaining period from 2016 to 2023.
@@ 5.2.2 Methods and implementation details
We assess the $m=9$ different forecasting models by the CRPS and construct SMCSs for the weak hypothesis as proposed in Section 3.4. By predictability of the forecasts, the CRPS differences are conditionally bounded. More precisely, for any predictive cdfs $F_{1}$ and $F_{2}$, we have
$\lvert\textrm{CRPS}(F_{1},y)-\textrm{CRPS}(F_{2},y)\rvert\leq\max\{\lvert\textrm{CRPS}(F_{1},z)-\textrm{CRPS}(F_{2},z)\rvert\mid z\in\mathcal{X}\},\quad y\in\mathbb{R},$ (19)
[…]
For each station, we thus compute the crossing points of all issued forecasts at all days in order to obtain predictable bounds of the CRPS differences by (19). In a second step, we use the transformation of Section 3.5 to obtain uniformly bounded score differences. That is, $\lvert d_{ij,t}^{k}\rvert\leq C/2$ for the resulting CRPS differences at station $k\leq K$, for $C=2$, $i,j\leq m$ and $t\in\mathbb{N}$. We let $\lambda=(2C)^{-1}=1/4$, compare Section 3.5, and choose $\alpha=0.1$ as our confidence level. In order to draw inference over all stations simultaneously, we consider the average scaled CRPS differences ${d}_{ij,t}=\sum_{k=1}^{K}{d}^{k}_{ij,t}/K$, for $i,j\leq 9$ and $t\in\mathbb{N}$.
@@ 5.2.3 Results
First, we take a look at the SMCS with respect to the loss differences averaged over all stations, from which we draw conclusions regarding the performance of postprocessing methods that align with the results in Schulz and Lerch (2022). The left panel of Figure 5 displays the resulting SMCS and shows that the two neural network-based approaches DRN and BQN, which performed best in the former study, are included in the SMCS for the entire time period. The other models fall out of the SMCS over time, where the order coincides with that found in the former study, e.g., the underlying ensemble is omitted first. In this case, where we average the performances over all stations, the SMCS behaves as for the stronger hypotheses, that is, once we eliminate a model, we do not include it anymore.
@@
Second, we consider the SMCS station-wise. The right panel of Figure 5 gives a summary, as it shows the number of stations for which a forecasting model is included in the corresponding SMCS. We find that the EPS is excluded from most SMCSs already within the first year, that the basic techniques EMOS, MBM and IDR are also omitted relatively fast, and that the machine learning approaches EMOS-GB and QRF fall out of SMCSs successively over time, until early 2021 when the last NWP model update occurred. Interestingly, the number of stations that include the HEN network approach increases after the last NWP model update, a behavior only detectable using the weak hypothesis. Finally, the BQN and DRN approach are included at almost all stations over the entire period, only with DRN being excluded over the last year at some stations.
Refer to captionRefer to caption
@@
Figure 5: SMCS averaged over the loss differences over all stations (left) and the number of stations where the method is included in the SMCS for each method dependent on the time (right). The vertical black lines indicate the three major NWP model updates.
@@
Taking a closer look at the individual stations, we observe different kinds of local behavior that is not visible when averaging over the loss differences. According to the findings, we broadly divide the stations into three categories: Regular (56 stations), dynamic (55) and fluctuating stations (55). Figure 6 shows one exemplary station for each of the three categories as well as the mean size of the station-wise SMCSs over time. The SMCS of a regular station behaves (almost) equally as for the stronger hypotheses, which is also seen in the plot of the mean SMCS size over the time period. The dynamic stations show a different pattern. As for the regular stations, some of the inferior models are excluded after a certain time period. However, in contrast to the regular stations, some methods that have been eliminated before are now included again coming closer to the end of the time period. This behavior becomes apparent when looking at the mean size of the SMCSs, which increases after the last update of the underlying NWP model that systematically changes the predictive performance of the methods. These cases are interesting, since any test which would be performed just at the end of the evaluation period (or at some given time point in the middle) would probably not detect the deficiencies of these models. Consider for example the SMCS at station 10044, Leuchtturm Kiel, displayed at the topright of Figure 6, where EMOS, IDR and HEN have been excluded from the SMCS for long periods and have become competitive again later on. Even though these three models perform not as strongly as the benchmarks DRN and BQN, the SMCS indicates that they are still better than the remaining models for this station. At last, we have the fluctuating stations, where at least one model is repeatedly excluded and included from the SMCS. It shows a kind of ”borderline”-behaviour of the method and indicates that it is exactly at the border of being worse than the other models. As for the other two groups, this behavior is reflected in the mean size of the SMCS, which is more oscillating for the fluctuating stations than the other two categories. In all cases, the sequential nature of the SMCS yields a deeper insight into the performances of the methods than tests for predictive ability which are issued at single time points only.
Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 6: SMCSs for one exemplary station in each of the three categories (topleft: normal; topright: dynamic; bottomleft: fluctuating) and the mean size of the SMCS averaged over the stations of each category dependent on the time (bottomright). Figure 8 in Appendix G shows the spatial distribution of all station categories over Germany.
@@ 6 Discussion
We introduced sequential model confidence sets and provided techniques to construct them with respect to three important notions of forecast superiority. Sequential model confidence sets address a highly relevant problem and allow to continuously monitor the performances of some given statistical models. They incorporate the uncertainty of the undergoing selection by being flexible in size, rely on minimal distributional assumptions, and achieve safe anytime-valid guarantees.
Concerning future work, there are various possible directions to investigate and extend the methods developed so far: Firstly, our construction is limited to (conditionally) bounded score differences. Important loss functions such as the quadratic score or the logarithmic score do not satisfy this requirement, and it would be useful to study e-processes and SMCSs for these more general loss functions, compare to Remark 3. Secondly, the simulation study shows that our methods are conservative. Since we have to adjust our processes for multiple testing, we may lose a considerable amount of evidence. More work is needed to reduce the gap in power between the raw and the adjusted e-processes, possibly by adaptive weighting schemes of the e-processes. Alternatively, one could study sequential model confidence sets with time-uniform FDR control as discussed in Appendix D. However, our impression is that there is not much gain in power by replacing the FWER with the less restrictive FDR. Finally, it would be interesting to examine how our results extend to important information criteria, and to apply sequential model confidence sets for sequential model selection.
@@
Code in R (R Core Team, 2024) for the case studies and replication material for the simulations are available at https://github.com/GGavrilos/SMCS.
@@ Acknowledgments
We thank Lutz Dümbgen for valuable discussions and inputs. Sebastian Arnold and Johanna Ziegel gratefully acknowledge financial support from the Swiss National Science Foundation. Computations for the simulations and parts of the wind gust case study have been performed on UBELIX (https://ubelix.unibe.ch/), the HPC cluster of the University of Bern. Benedikt Schulz gratefully acknowledges funding within the project C5 “Dynamical feature-based ensemble postprocessing of wind gusts within European winter storms” of the Transregional Collaborative Research Center SFB/TRR 165 “Waves to Weather” funded by the German Research Foundation (DFG). We thank Cristina Primo, Reinhold Hess and Sebastian Trepte for providing the forecast and observation data, and Robert Redl for assistance in data handling.
Appendix A Proper scoring rules and consistent scoring functions
@@
In this section, we review some important concepts for forecast evaluation. We assume that we are interested in a real-valued unknown quantity and let $\mathcal{P}(\mathbb{R})$ be the family of all probability distributions on $\mathbb{R}$.
@@
A statistical functional is a map $T:\mathcal{P}(\mathbb{R})\to 2^{\mathbb{R}}$. Point-valued forecasts for functionals should be compared using consistent scoring functions (Gneiting, 2011). A measurable map $S:\mathbb{R}\times\mathbb{R}\rightarrow\bar{\mathbb{R}}$ is a consistent scoring function for the functional $T$ relative to the class $\mathcal{P}\subseteq\mathcal{P}(\mathbb{R})$ if for all $x\in\mathbb{R}$, $F\in\mathcal{P}$, the integral $\int S(x,y)\,\mathrm{d}F(y)$ exists and
$\int S(t,y)\,\mathrm{d}F(y)\leq\int S(x,y)\,\mathrm{d}F(y),\quad\textrm{for all }t\in T(F).$
The Brier score (or quadratic score) $\textrm{BS}(x,y)=(x-y)^{2}$, $x,y\in\mathbb{R}$, and the quantile score
$\textrm{QS}_{\tau}(x,y)=(\mathbbm{1}{\{y\leq x\}}-\tau)(x-y),\quad x,y\in\mathbb{R},$ (20)
are main examples of consistent scoring functions for the mean functional, and for the quantile at level $\tau\in(0,1)$, respectively. More generally, Ehm et al. (2016) show that a scoring function $S$ is consistent for the quantile functional at level $\tau\in(0,1)$ if and only if it is of the form
$S_{\tau}(x,y)=(\mathbbm{1}{\{y\leq x\}}-\tau)\big{(}g(x)-g(y)\big{)},\quad x,y\in\mathbb{R},$
for some increasing function $g:\mathbb{R}\to\mathbb{R}$, where we obtain the quantile score at (20) for $g$ equals the identity, and where the result extends for the mean and the expectile.
Probabilistic forecasts incorporate the intrinsic uncertainty of future events and comprise of a probability measure on the possible outcomes. Probabilistic forecasts should be compared and evaluated using proper scoring rules (Gneiting and Raftery, 2007). A proper scoring rule is a function ${\rm S}:\mathcal{P}\times\mathbb{R}\mapsto\bar{\mathbb{R}}$ for some class $\mathcal{P}\subseteq\mathcal{P}(\mathbb{R})$ such that ${\rm S}(F,\cdot)$ is measurable for any $F\in\mathcal{P}$, the integral $\int{\rm S}(G,y)F(\,\mathrm{d}y)$ exists, and
$\int{\rm S}(F,y)\,\mathrm{d}F(y)\leq\int{\rm S}(G,y)\,\mathrm{d}F(y),\quad\textrm{for all }F,G\in\mathcal{P}.$
@@
That is, for a proper scoring rule ${\rm S}$, the expected score $\mathbb{E}_{Y\sim F}{\rm S}(G,Y)$ is minimized over all distributional forecasts $G\in\mathcal{P}$, if we correctly forecast the true distribution of the random variable $Y$. The most commonly used proper scoring rule for the evaluation of real-valued outcomes is the continuous ranked probability score (CRPS; Matheson and Winkler (1976)), which is defined for all $F\in\mathcal{P}(\mathbb{R})$ and $y\in\mathbb{R}$ as
$\textrm{CRPS}(F,y)=\int(F(x)-\mathbbm{1}\{y\leq x\})^{2}\,\mathrm{d}x.$ (21)
@@
The CRPS is a continuous extension of the Brier score for binary events and is popular across application areas and methodological communities; see, e.g., Gneiting et al. (2005), Hothorn et al. (2014) and Henzi et al. (2021).
Appendix B Conditionally bounded CRPS and quantile score differences

### v3

@@ 5.2 Postprocessing of wind gust predictions
Weather services use numerical weather prediction (NWP) models for predicting the future state of the atmosphere. These NWP models quantify forecast uncertainty using ensemble predictions, where each ensemble member represents a different future scenario. However, such ensemble forecasts are typically subject to systematic biases and dispersion errors. Hence, the forecasts ought to be statistically postprocessed to generate accurate and reliable predictions (Gneiting and Raftery, 2005; Vannitsem et al., 2018).
Schulz and Lerch (2022) present a comprehensive comparison of multiple approaches for statistical postprocessing of ensemble forecasts for wind gusts. Here, we will build on their case study and compare the different forecasting models for probabilistic wind gust prediction by using SMCSs under the weak hypothesis.
@@ 5.2.1 Data
We use the data set in Primo et al. (2024), which is an extension of the original data in Schulz and Lerch (2022). The forecast data is based on an operational NWP model from the German weather service (Deutscher Wetterdienst, DWD) in the period from 08/12/2010 to 30/06/2023 with corresponding observations at 174 geographically diverse weather stations in Germany. The forecasts are initialized at 00 UTC and have a forecast horizon of up to 21 hours. Here, we consider only those forecasts with a lead time of 18h referring to 18 UTC and the $K=166$ stations with at least 2,600 observations.
Next to the ensemble forecasts (EPS), we include the eight postprocessing methods applied by Schulz and Lerch (2022). Statistical postprocessing methods are typically distributional regression models that use ensemble predictions from the NWP model as input data. The methods in we compare can be divided into three groups: Basic techniques that only make use of the underlying wind gust ensemble predictions (EMOS, MBM, IDR), established machine learning methods for postprocessing that incorporate additional covariates (EMOS-GB, QRF), and neural network-based approaches (DRN, BQN, HEN). While the first two groups fit a separate model for each station, the neural networks estimate one locally adaptive model for all stations. Further, the models differ in the resulting forecast distributions, which range from parametric (EMOS, EMOS-GB, DRN) to semi- and non-parametric types (all others). The postprocessing models are trained on the period from 2010 to 2015 using the same configuration as in Schulz and Lerch (2022) and evaluated on the remaining period from 2016 to 2023.
@@ 5.2.2 Methods and implementation details
We assess the $m=9$ different forecasting models with the CRPS and construct SMCSs for the weak hypothesis as proposed in Section 3.3. Since the forecasts are predictable, the CRPS differences are conditionally bounded, see Supplement B. We use the transformation we discussed in Section 3.4 to obtain uniformly bounded score differences. That is, $\lvert d_{ij,t}^{k}\rvert\leq c/2$ for the resulting CRPS differences at station $k\leq K$, for $c=2$, $i,j\leq m$, $t\in\mathbb{N}$. We set $\lambda_{ij,t}=(2c)^{-1}=1/4$ and $\alpha=0.1$. For inference on all stations simultaneously, we consider the average scaled CRPS differences ${d}_{ij,t}=\sum_{k=1}^{K}{d}^{k}_{ij,t}/K$, for $i,j\leq 9$ and $t\in\mathbb{N}$.
@@ 5.2.3 Results
First, for the SMCS with respect to the loss differences averaged over all stations, our conclusions regarding the performance of postprocessing methods align with the results in Schulz and Lerch (2022). The left panel of Figure 5 displays the SMCS and shows that the two neural network-based approaches, DRN and BQN, which performed best in the former study, are included in the SMCS for the entire time period. The other models fall out of the SMCS over time, where the order coincides with that found in the former study, e.g., the EPS is omitted first. In this case, where we average the performances over all stations, the SMCS behaves as for the stronger hypotheses, that is, once we eliminate a model, we do not include it anymore.
@@
Second, we consider the SMCS station-wise. The right panel of Figure 5 gives a summary, as it shows the number of stations for which a forecasting model is included in the corresponding SMCS. The EPS is excluded from most SMCSs already within the first year, the basic techniques EMOS, MBM and IDR are also excluded relatively fast, and that the machine learning approaches EMOS-GB and QRF fall out of the SMCSs successively over time, until early 2021 when the last NWP model update occurred. Interestingly, the number of stations that include the HEN network approach increases after the last NWP model update, a behavior only detectable using the weak hypothesis. Finally, the BQN and DRN approach are included at almost all stations over the entire period, only with DRN being excluded over the last year at some stations.
@@
Figure 5: SMCS of averaged loss differences over all stations (left) and the number of stations where a method is included in the SMCS dependent on time (right). The vertical black lines indicate the three major NWP model updates.
@@
Taking a closer look at the individual stations, we observe different kinds of local behavior that is not visible when averaging over the loss differences. We can broadly divide the stations into three categories: Regular (56 stations), dynamic (55) and fluctuating(55). Figure 6 shows one exemplary station for each of the three categories as well as the mean size of the station-wise SMCSs over time. The SMCS of a regular station behaves (almost) equally as for the stronger hypotheses, which is also seen in the plot of the mean SMCS size over the time period. The dynamic stations show a different pattern. As for the regular stations, some of the inferior models are excluded after a certain time period. However, in contrast to the regular stations, some methods that have been eliminated before are now included again, coming closer to the end of the time period. This behavior becomes apparent when looking at the mean size of the SMCSs, which increases after the last update of the underlying NWP model, which systematically changes the predictive performance of the methods. These cases are interesting, since any test that would be performed just at the end of the evaluation period (or at some given time point in the middle) may not detect the deficiencies of these models. Consider, for example, the SMCS at station 10044, Leuchtturm Kiel, displayed at the top right of Figure 6, where EMOS, IDR and HEN are excluded from the SMCS for long periods and become competitive again later on. Although these three models do not perform as strongly as DRN and BQN, the SMCS indicates that they are still better than the remaining models for this station. Finally, we have the fluctuating stations, where at least one model is repeatedly excluded and included from the SMCS. As for the other two groups, this behavior is reflected in the mean size of the SMCS, which is oscillating more for the fluctuating stations than the other two categories. In all cases, the sequential nature of the SMCS yields a deeper insight into the performance of the methods than tests for predictive ability which are issued at single time points only.
@@
Figure 6: SMCSs for one exemplary station in each of the three categories (top left: normal; top right: dynamic; bottom left: fluctuating) and the mean size of the SMCS averaged over the stations of each category dependent on the time (bottom right).
@@ 6 Discussion
We introduce sequential model confidence sets and provide methodology to construct them with respect to three important notions of forecast superiority. SMCSs allow to continuously monitor the performance of some given statistical models. They incorporate the uncertainty in choosing models by being flexible in size, rely on minimal distributional assumptions, and achieve anytime-valid coverage guarantees.
We have provided coverage guarantees of all superior models with high probability uniformly over time. Other coverage guarantees could be of interest but require somewhat different approaches. The literature on inference on the argmin of a vector of means typically focuses on the guarantee that each superior model is covered with high probability. Future work could consider time-uniform extensions of such guarantees. Furthermore, one could aim for a time-uniform bound of the expected fraction of wrongly rejected models by the total number of rejected models. This can be achieved by time-uniform control of the false discovery rate and we give first results in this direction in Supplement D.
We have mainly focused on (conditionally) bounded score differences. However, important loss functions such as the squared error or the logarithmic score do not satisfy this requirement. Nevertheless, our general theory is not limited to bounded score differences. We have provided a simulation study for SMCSs with the squared error loss in Supplement F.1. Nonetheless, further research is warranted to extend the understanding and implementation of SMCSs to encompass more general loss functions. Finally, it would be interesting to examine how our results extend to important information criteria, and to apply sequential model confidence sets for sequential model selection.
@@
Code in R for the case studies and replication material for the simulations are available at https://github.com/GGavrilos/SMCS.
@@ Acknowledgments
The comments of three anonymous referees significantly improved the quality of the paper. We thank Lutz Dümbgen for valuable discussions and inputs. Sebastian Arnold and Johanna Ziegel gratefully acknowledge financial support from the Swiss National Science Foundation. Computations for the simulations and parts of the wind gust case study have been performed on UBELIX (https://ubelix.unibe.ch/), the HPC cluster of the University of Bern. Benedikt Schulz gratefully acknowledges funding within the project C5 “Dynamical feature-based ensemble postprocessing of wind gusts within European winter storms” of the Transregional Collaborative Research Center SFB/TRR 165 “Waves to Weather” funded by the German Research Foundation (DFG). We thank Cristina Primo, Reinhold Hess and Sebastian Trepte for providing the forecast and observation data, and Robert Redl for assistance in data handling.
Supplementary material
Appendix A Scoring rules and scoring functions
@@
In this section, we review some important concepts for forecast evaluation. We assume that the unknown quantity of interest maps into some measurable space $\mathcal{X}$ and let $\mathcal{P}(\mathcal{X})$ be the family of all probability distributions on $\mathcal{X}$.
@@
A statistical functional is a map $T:\mathcal{X}\to 2^{\mathcal{X}}$. Point-valued forecasts for functionals should be compared using consistent scoring functions (Gneiting, 2011). A measurable map $S:\mathcal{X}^{2}\rightarrow\bar{\mathbb{R}}$ is a consistent scoring function for the functional $T$ relative to the class $\mathcal{P}\subseteq\mathcal{P}(\mathcal{X})$ if for all $x\in\mathcal{X}$, $F\in\mathcal{P}$, the integral $\int S(x,y)\,\mathrm{d}F(y)$ exists and, for all $t\in T(F)$, $\int S(t,y)\,\mathrm{d}F(y)\leq\int S(x,y)\,\mathrm{d}F(y)$. For real-valued outcomes, that is $\mathcal{X}=\mathbb{R}$, the Brier score (or quadratic score) $\textrm{BS}(x,y)=(x-y)^{2}$, and the quantile score $\textrm{QS}_{\tau}(x,y)=(\mathbbm{1}{\{y\leq x\}}-\tau)(x-y)$ are main examples of consistent scoring functions for the mean functional, and for the quantile at level $\tau\in(0,1)$, respectively. A scoring function $S$ is consistent for the quantile functional at level $\tau\in(0,1)$ if and only if it is of the form $S_{\tau}(x,y)=(\mathbbm{1}{\{y\leq x\}}-\tau)\big{(}g(x)-g(y)\big{)},x,y\in\mathbb{R},$ for some increasing function $g:\mathbb{R}\to\mathbb{R}$. Similar characterization result for the mean and expectiles can be found in Gneiting (2011) alongside historical references for them.
@@
Probabilistic forecasts quantify the uncertainty of the future event and are specified as a probability measure over all possible outcomes. Probabilistic forecasts should be compared and evaluated using proper scoring rules (Gneiting and Raftery, 2007; Waghmare and Ziegel, 2025). A proper scoring rule is a function ${\rm S}:\mathcal{P}\times\mathcal{X}\mapsto\bar{\mathbb{R}}$ for some class $\mathcal{P}\subseteq\mathcal{P}(\mathcal{X})$ such that ${\rm S}(F,\cdot)$ is measurable for any $F\in\mathcal{P}$, the integral $\int{\rm S}(G,y)\,\mathrm{d}F(y)$ exists, and for all $F,G\in\mathcal{P}$, $\int{\rm S}(F,y)\,\mathrm{d}F(y)\leq\int{\rm S}(G,y)\,\mathrm{d}F(y)$. That is, for a proper scoring rule ${\rm S}$, the expected score $\mathbb{E}_{Y\sim F}{\rm S}(G,Y)$ is minimized with respect to all distributional forecasts $G\in\mathcal{P}$, if we correctly forecast the true distribution $F$ of the random variable $Y$. The most commonly used proper scoring rules for the evaluation of real-valued outcomes are the logarithmic score, $S(F,y)=-\log f(y)$, where $f$ is the density of $F$, and the continuous ranked probability score (CRPS; Matheson and Winkler (1976)), which is defined for all $F\in\mathcal{P}(\mathbb{R})$ with finite first moment and $y\in\mathbb{R}$ as
$\textrm{CRPS}(F,y)=\int(F(x)-\mathbbm{1}\{y\leq x\})^{2}\,\mathrm{d}x.$
@@
The CRPS is a continuous extension of the Brier score for binary events and is popular across application areas and methodological communities; see, e.g., Gneiting et al. (2005), Henzi et al. (2021), and Waghmare and Ziegel (2025) for a comprehensive recent review.
Propriety of scoring rules is maintained under normalization by predictable quantities in the following sense: Let $(\mathcal{F}_{t})_{t\in\mathbb{N}}$ be some filtration and $S$ be a proper scoring rule. Then, $\tilde{S}_{t}=S/c_{t}$ is conditionally proper for any predictable $(c_{t})_{t\in\mathbb{N}}\subseteq(0,\infty)$, that is, for any predictable sequence of probabilistic forecasts $(G_{t})_{t\in\mathbb{N}}$, and any adapted $(Y_{t})_{t\in\mathbb{N}}$ with conditional distribution $F_{t}$ given $\mathcal{F}_{t-1}$, we have
$\mathbb{E}[{\tilde{S}_{t}(F_{t},Y_{t})}\mid\mathcal{F}_{t-1}]=c_{t}^{-1}\mathbb{E}\left[S(F_{t},Y_{t})\mid\mathcal{F}_{t-1}\right]\leq c_{t}^{-1}\mathbb{E}\left[S(G_{t},Y_{t})\mid\mathcal{F}_{t-1}\right]=\mathbb{E}[{\tilde{S}_{t}(G_{t},Y_{t})}\mid\mathcal{F}_{t-1}],$
where the inequality follows by propriety of $S$ applied conditionally on $\mathcal{F}_{t-1}$.
Appendix B Conditionally bounded score differences